1. Say, Mr. X runs on a plain land and jumps from a spot in forward direction. Recognize and




describe his jumping motion through an equation and make an argument for your answer.

1
Expert's answer
2021-11-01T12:53:27-0400

"\text{Let at the moment of the jump mr.X has the following speeds:}"

"v_x - \text{instantaneous speed of movement horizontally }"

"v_y - \text{instantaneous vertical speed}"

"\text{Then the coordinates mr.X will be:}"

"x = v_xt(1)"

"y = v_yt -\frac{gt^2}{2}(2)"

"\text{The equations (1) and (2) coincide with the parametric}"

"\text{ equations of a body thrown at an angle to the horizon.}"

"\text{Where:}"

"v=\sqrt{v_x^2+v_y^2} \text{ -body speed}"

"\tan \alpha= \frac{v_y}{v_x};\alpha-\text{throw angle}"